OpenStudy (loser66):

Please, check my stuff, I don't know what is wrong!! Find all solutions of 102x +1001 y = 1

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

So, did you learn about the "zero's" of an equation yet or no??

OpenStudy (loser66):

1001 = (8) 120 + 41, right?

OpenStudy (loser66):

Hence 41 = 1001 -(8) 120

OpenStudy (loser66):

120 = (2)41 +38 , hence 38 = 120 - (2)41

OpenStudy (loser66):

41 = 38 + 3 , hence 3 = 41 -38 and 38 = (12)3 +2 , hence 2 = 38 -(12)3 3 = 2+ 1 1 = 3-2

OpenStudy (loser66):

Am I right so far?

OpenStudy (loser66):

Now, backward 1 = 3-2 = 3 -(38 -(12)3 ) = (13)3 -38

OpenStudy (anonymous):

For the equation 102x + 1001y = 1, you know that the gcd of 102 and 1001 is 1. Therefore, there is some linear combination of 102 and 1001 that adds to 1. Use the Euclidean Algorithm to find the coefficients m, n so that 102 n + 1001 m = 1. Once you find that m and n, note that you can generate an infinite set of solutions by noting: 102 (n + 1001) + 1001 (m - 102) = 1 is always a true statement if n and m is a solution to 102 n + 1001 m = 1. Also, 102 (n + 2*1001) + 1001 (m - 2*102) = 1 is true. 102 (n + 1001k) + 1001 (m - 102)k = 1 for any integer k is true as well. So we have an infinite set of solutions based on our initial solution given by the Euclidean algorithm to find the gcd of two numbers.

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

I need a concrete solution, i.e. numbers in answer, not m, k, or n

OpenStudy (loser66):

and I am on the way, just don't know where I messed up. So that I need someone check my steps

OpenStudy (loser66):

1= (13) 3 -38 = (13 )(41-38) -38 = (13) 41 -(14)38 1= (13) 41 -(14) (120 -(2)41) = (13)41 -(14)120 +(28) 41= (41) 41 -(14)120

OpenStudy (loser66):

1= (41)(1001-(8)`120)-(14)120 = (41) 1001-(328)120 -(14)120 = (41) 1001 -(342) 120

OpenStudy (loser66):

ok, so far so good. Hence x = 41 and y = -342, and that is only one solution, right?

OpenStudy (loser66):

oh, no, still have other. Since gcd (1001,120) = 1, and 1|1, hence the solutions should be \(X = x + 1001t \rightarrow X = 41 + 1001 t\) and \(Y= y - 102 t \rightarrow Y = -342 -102 t\) for all t in \(\mathbb Z\)

OpenStudy (loser66):

But I don't know where I messed up. I don't get the correct one.

OpenStudy (loser66):

I mean, if t = 1, then X = 41 + 1001 = 1042 Y = -342 -102 = -444 Plug back to original one. 102 X +1001Y = 102 (1042) + 1001(-444) = 106284-444444=-338160 while I am supposed to get 1 back!!!

ganeshie8 (ganeshie8):

is that a linear diophantine eqn ?

OpenStudy (loser66):

Yes, it is @ganeshie8

ganeshie8 (ganeshie8):

what is your one particular solution ?

OpenStudy (loser66):

oh, I see where I messed up, at the very first step 1001 = (9)102 +83

OpenStudy (loser66):

Not 8(120) +41 OMG

OpenStudy (loser66):

@ganeshie8 Let me post my first try with it, and still get the wrong answer :(

OpenStudy (loser66):

Yes

ganeshie8 (ganeshie8):

Forward : ``` 1001 - (102 x 9) = 83 102 - (83 x 1) = 19 83 - (19 x 4) = 7 19 - (7 x 2) = 5 7 - (5 x 1) = 2 5 - (2 x 2) = 1 ```

OpenStudy (loser66):

Ok, got this. I follow you,

OpenStudy (loser66):

next, please

ganeshie8 (ganeshie8):

Next, start from last step

ganeshie8 (ganeshie8):

your goal is to express `gcd` as a linear combination of 1001 and 102

OpenStudy (loser66):

Yes, I know

ganeshie8 (ganeshie8):

Forward : ``` 1001 - (102 x 9) = 83 102 - (83 x 1) = 19 83 - (19 x 4) = 7 19 - (7 x 2) = 5 7 - (5 x 1) = 2 5 - (2 x 2) = 1 ``` Backward : ``` 1 = 5 - 2*2 = 5 - 2(7-5*1) = 5(1+2) + 7(-2) = 5(3) + 7(-2) ```

ganeshie8 (ganeshie8):

fine with that step ?

OpenStudy (loser66):

i got (422)102 -(43) 1001 =1 . Finally!!

OpenStudy (loser66):

Hence \(x_0 = 422; y_0 = -43\) right?

ganeshie8 (ganeshie8):

good, so (422, -43) is one particular solution to the eqn 102x + 1001y = 1

ganeshie8 (ganeshie8):

next find the null solution

OpenStudy (loser66):

What is null solution? I know general solution only That is \(X = x_0 + 1001 t\) and \(Y= y_0 - 102 t\)

ganeshie8 (ganeshie8):

t*(1001, -102) is called the null solution

ganeshie8 (ganeshie8):

particular solution : (422, -43) null solution : t*(1001, -102) complete solution = particular + null = (422, -43) + t*(1001, -102)

ganeshie8 (ganeshie8):

if you see, the points `t*(1001, -102)` are the solutions to the equaiton 102x + 1001y = 0. thats the reason it is called a null solution

OpenStudy (loser66):

Yes, I saw it. Thank you so much. It is not hard, right? just easy make mistake and it drove me crazy when I can't find where the mistake is. :(

ganeshie8 (ganeshie8):

it's not hard as it is a very familairnew pattern, you must have seen this kind of constructing solutions in linear algebra and differential equations

ganeshie8 (ganeshie8):

you find any one solution, find the null solution (this is almost always trivial) write out the complete solution

OpenStudy (loser66):

Yes. Thank you very much

jimthompson5910 (jim_thompson5910):

Here's a slightly different way to do it. It's basically the extended euclidean algorithm, but for me, it helps to have variables thrown in 1001 = 9*102 + 83 -----> 83 = 1001 - 9*102 102 = 1*83 + 19 -----> 19 = 102 - 1*83 83 = 4*19 + 7 -----> 7 = 83 - 4*19 19 = 2*7 + 5 -----> 5 = 19 - 2*7 7 = 1*5 + 2 -----> 2 = 7 - 1*5 5 = 2*2 + 1 -----> 1 = 5 - 2*2 let x = 1001 and y = 102 83 = 1001 - 9*102 83 = x - 9*y ------------------ 19 = 102 - 1*83 19 = y - 1*(x - 9*y) 19 = 10y - x ------------------ 7 = 83 - 4*19 7 = (x - 9*y) - 4*(10y - x) 7 = 5x-49y ------------------- 5 = 19 - 2*7 5 = (10y - x) - 2*(5x-49y) 5 = 108y-11x ------------------- 2 = 7 - 1*5 2 = (5x-49y) - 1*(108y-11x) 2 = 16x-157y ------------------- 1 = 5 - 2*2 1 = (108y-11x) - 2*(16x-157y) 1 = 422y-43x -43x + 422y = 1 -43(1001) + 422(102) = 1 1001*(-43) + 102*(422) = 1

jimthompson5910 (jim_thompson5910):

or you can use the method described here https://en.wikibooks.org/wiki/Discrete_Mathematics/Number_theory#The_extended_Euclidean_algorithm to generate this table \[ \begin{array}{ccccc} i & q & r & u & v\\ -1 & X & 1001 & 0 & 1\\ 0 & X & 102 & 1 & 0\\ 1 & 9 & 83 & -9 & 1\\ 2 & 1 & 19 & 10 & -1\\ 3 & 4 & 7 & -49 & 5\\ 4 & 2 & 5 & 108 & -11\\ 5 & 1 & 2 & -157 & 16\\ 6 & 2 & 1 & \color{red}{422} & \color{red}{-43}\\ 7 & 2 & 0 & X & X\\ \end{array} \] 102u + 1001v = 1 102(422) + 1001(-43) = 1